OpenStudy (anonymous):

If Jessie is 24 years younger than her mother and if the sum of their ages is 84, how old is Jessie?

OpenStudy (anonymous):

J M=J+24 (J+24)+J=84

OpenStudy (anonymous):

2J+24=84

OpenStudy (anonymous):

J is jessie's age while M is her mom's age

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

2J=84-24 2J=60 J=30